OpenStudy (anonymous):

Solve the equation: 2x+1/3 - x-1/4 = 0 Please show the steps !!!

OpenStudy (anonymous):

Is this your question: \[\frac{2x+1}{3}-\frac{x-1}{4}=0\]

OpenStudy (anonymous):

Yes that's the problem

OpenStudy (anonymous):

If so, multiply all terms by the GCF and then use the distributive property to solve.

OpenStudy (anonymous):

And the distributive property is:  Does this help?